The level of activity at the volcano remained low today. There were 3 volcano-tectonic earthquakes and 16 rockfalls.
COSPEC measurements made from the MV Shamrock II yesterday gave a sulphur dioxide flux of 1713 tonnes per day. This elevated flux may have been related to the enhanced steam venting noticed yesterday.
Over the last few days GPS surveys have been carried out using the following sites: Brodericks, Whites, Long Ground, Windy Hill, Lees Yard, Galways Estate and the two base stations at Harris and Hermitage. The results processed so far show that there has been no significant change in the pattern of deformation around the volcano.
The rain showers today have kept the dust levels low at all sites.
Residents should continue listening to Radio Montserrat in case there are sudden developments in the state of the volcano.